# AssemAI: Interpretable Image-Based Anomaly Detection for Manufacturing Pipelines

### v1

@@ Abstract
Anomaly detection in manufacturing pipelines remains a critical challenge, intensified by the complexity and variability of industrial environments. This paper introduces AssemAI, an interpretable image-based anomaly detection system tailored for smart manufacturing pipelines. Our primary contributions include the creation of a tailored image dataset and the development of a custom object detection model, YOLO-FF, designed explicitly for anomaly detection in manufacturing assembly environments. Utilizing the preprocessed image dataset derived from an industry-focused rocket assembly pipeline, we address the challenge of imbalanced image data and demonstrate the importance of image-based methods in anomaly detection. The proposed approach leverages domain knowledge in data preparation, model development and reasoning. We compare our method against several baselines, including simple CNN and custom Visual Transformer (ViT) models, showcasing the effectiveness of our custom data preparation and pretrained CNN integration. Additionally, we incorporate explainability techniques at both user and model levels, utilizing ontology for user-friendly explanations and SCORE-CAM for in-depth feature and model analysis. Finally, the model was also deployed in a real-time setting. Our results include ablation studies on the baselines, providing a comprehensive evaluation of the proposed system. This work highlights the broader impact of advanced image-based anomaly detection in enhancing the reliability and efficiency of smart manufacturing processes.
@@
Anomaly Detection, Object Detection, Smart Manufacturing, Interpretability, Image Processing
@@ I Introduction
Anomaly detection is an essential methodology in manufacturing to identify deviations from normal production processes, which can indicate potential issues such as equipment failures, defects, or inefficiencies. Anomaly detection helps ensure product quality, reduce downtime and optimize production processes [3] [4]. By monitoring various process parameters and product characteristics, manufacturers can detect anomalies early and prevent costly production disruptions [5].
@@
Image data is increasingly being leveraged in manufacturing systems due to advancements in computer vision and the availability of high-resolution cameras. In modern manufacturing facilities, image data is utilized for a wide range of applications, including quality inspection [6, 7], predictive maintenance [8, 9], process monitoring [10] and process optimization [11]. The integration of image data allows for real-time monitoring and anomaly detection, significantly improving the ability to identify defects and streamline production processes. Specifically, image-based anomaly detection has gained prominence, leveraging visual data to identify defects or irregularities in products. Automated visual inspection systems use image data to detect surface defects, assembly errors and deviations from design specifications[12], providing real-time feedback essential for maintaining high production standards and minimizing defects[4].
@@
Despite the advancements in image-based anomaly detection, several challenges persist. (i) High variability in appearance between normal and anomalous conditions, which can make it difficult for models to differentiate subtle defects from acceptable variations. (ii) Requirement for large and diverse datasets for training, which can be resource-intensive to collect and annotate [5, 13]. (iii) Challenges with generalization, where a model trained on images from one manufacturing environment may not perform well in different settings due to variations in lighting, camera angles, or product types [14]. (iv) Difficulty in interpreting detection results, affecting the ability to provide actionable insights for process improvement [15].
Figure 1: Overall Architecture. The figure includes the pipeline for AssemAI, beginning with dataset preparation involving filtering and cropping images, followed by zero-shot object detection using two approaches (a custom model and YOLO-FF), and culminating in object detection modeling with CNN, Custom-ViT, Pretrained-ViT, and Efficient-NET. The detection output is then explained using SCORE-CAM for model-level explanations and integrated with process ontology for user-level explanations.
To address these challenges, this work presents the following contributions:
@@
  • A novel anomaly detection model utilizing the pre-trained EfficientNet architecture [16], fine-tuned to a new dataset for improved anomaly detection in assembly processes.
  • A new image dataset for anomaly detection in manufacturing, curated using a custom-trained YOLOv9 model named YOLO-FF (You Only Look Once-Future Factories).
@@
  • Integration of advanced explainability techniques, incorporating an ontology-based method for user-level explanations and SCORE-CAM for detailed feature importance and model analysis.
  • The image dataset and code to reproduce the results and additional experiments are available at this link.
@@ II-A Object Detection Models
Object detection has evolved with various models, improving both accuracy and efficiency. Girshick et al. [17] introduced R-CNN, which applies CNNs to region proposals for object classification and bounding box refinement. Faster R-CNN [18] built on R-CNN by integrating a Region Proposal Network (RPN) for faster detection. The YOLO models, from YOLOv1 to YOLOv9, transformed object detection by formulating it as a single regression problem to predict bounding boxes and class probabilities directly [19]. ShuffleNet [20] and MobileNet [21] are lightweight models designed for efficiency on mobile devices. SqueezeNet [22] aimed to reduce model size while maintaining accuracy with its Fire module. Lastly, the Swin Transformer [23] introduced hierarchical transformers with shifted windows, setting new benchmarks for object detection and vision tasks. Despite these advancements, limitations such as the need for extensive labeled data and computational resources persist. Our work addresses these limitations by focusing on lightweight and efficient models tailored for anomaly detection in manufacturing settings, improving both detection speed and accuracy while maintaining interpretability.
@@ II-B Zero-Shot Object Detection
Zero-shot object detection (ZSD) recognizes objects without labeled training data but faces challenges like semantic noise and class imbalance. Foundational methods by Bansal et al. [24] and Rahman et al. [25] struggled with noise. Gupta et al. [26] and Zheng et al. [27] improved ZSD with symmetric mapping and cascade stages, but issues persisted. Hayat et al. [28] and Li et al. [29] used ResNet, KNN, and contextual information but faced noise and ambiguity. Zhu et al. [30] and Hayat et al. [31] advanced ZSD with DELO and Generative-ZSD. Liu et al. [32] and Li et al. [33] proposed contrastive learning and a semantics-aware framework. Our work focuses on specialized domains for precise anomaly detection.
@@ II-C Anomaly detection
Anomaly detection has seen significant advancements across diverse domains like networking [34], smart agriculture [35], healthcare [36], foundational [35, 37, 38, 39], manufacturing [40, 41, 42, 43, 44, 45]. Kim et al. introduced NetViewer [34], a network measurement approach that detects, identifies and visualizes attacks and anomalous traffic in real-time by passively monitoring packet headers and representing them as frames or images. An anomaly detection method using Generative Adversarial Networks (GAN), which searches for good representations of samples in the generator’s latent space and deems samples without such representations as anomalous, was proposed by Deecke et al. [37]. Minhas et al, in their work addresses visual defect assessment for anomaly detection in tasks [38], like surface inspection by proposing a network-based deep transfer learning method using CNNs, hypothesizing that deeper networks will outperform traditional single class SVMs.
II-D Anomaly detection in manufacturing
Significant research has investigated image-based anomaly detection in manufacturing. Haselmann et al. [39] present an unsupervised one-class learning method using a deep CNN for surface inspection, outperforming other methods on decorated plastic parts. Xie et al. [40] propose a uniform benchmark for assessing image anomaly detection (IAD) algorithms in industrial settings. Maggipinto et al. [41] use convolutional autoencoders for monitoring semiconductor manufacturing, enhancing effectiveness and scalability. Jiang et al. [42] introduce YOLOv3 for balanced datasets and Fast-AnoGAN for unbalanced datasets in industrial production. Tan et al. [43] use an encoder-decoder for anomaly detection in sequential sensor data. Kim et al. [44] propose a self-supervised method using Gramian angular field and StyleGAN for time-series data. Bougaham et al. [45] demonstrate a three-step deep learning approach for Printed Circuit Board Assembly (PCBA) images, achieving high accuracy. Despite advancements, publicly available datasets for assembly processes are scarce, and existing methods lack interpretability and explainability for domain-specific insights.
@@ III Proposed Methodology
Figure 1 shows the overall architecture of the proposed method. The AssemAI pipeline begins with dataset preparation, where images are filtered and cropped to focus on relevant features. This is followed by zero-shot object detection using two approaches: a custom model and YOLO-FF, which adapts zero-shot detection for anomaly detecti use cases. Subsequently, object detection modeling is performed using several architectures, including CNN, Custom-ViT, Pretrained-ViT, and Efficient-NET, to ensure robust anomaly detection. The detection output is then explained using SCORE-CAM for model-level explanations and integrated with process ontology for user-level explanations, providing both technical and domain-specific insights.
III-A Future Factories Multimodal Dataset
We use the Future Factories (FF) dataset [46] generated and made publicly available by the Future Factories (FF) lab operating at the McNair Aerospace Research Center at the University of South Carolina. The dataset consists of measurements from a simulation of a rocket assembly pipeline, which adheres to industrial standards in deploying actuators, control mechanisms and transducers. The dataset is comprised of two versions; Analog and Multi-modal dataset, for which we use the images included in the Multi-modal dataset for our purpose. Table I shows the statistics of the original images in the FF Multimodal dataset.
@@ III-A Future Factories Multimodal Dataset
III-B Image filtering
The rocket assembly process at the FF lab is divided into 21 distinct cycle states. Information about these cell cycle states was not directly available in the multi-modal dataset and had to be extracted from the analog dataset using a mapping function provided by domain experts. By calculating the Structural Similarity Index (SSIM) between the normal and anomalous images as shown in Figure 2, we observed that the images are structurally very similar across most cycle states. This indicates that the differences between the images are subtle and localized, which suggests that further analysis is needed to filter and figure out the region of interest. Among these cell cycle states, the rocket and its parts are visible only in two specific cycle states due to the spatial location of robots and other machinery and the location of the cameras and camera angles. We focus on filtering cycle four and a section of cycle nine based on domain knowledge and observational insights for enhanced image understanding. Table II and III show the filtered dataset and its anomaly statistics, respectively.
Figure 2: Structural Similarity between Normal and Anomalous Images
@@ III-C Image Cropping
Our dataset comprises images captured at various stages of the manufacturing cycle. Each cycle consists of defined time stamps where the toy rocket’s position remains consistent. Leveraging this domain knowledge, we have implemented a cropping strategy to remove the background from all images. This consistent positioning of the toy rockets across cycles enables us to crop the images accurately, ensuring that only the relevant parts of the images are retained.
The necessity for this cropping approach became evident after our initial attempts at anomaly detection without cropping yielded sub-optimal results. Specifically, our models performed poorly, with significant misclassification rates. We employed the explainability model SCORE-CAM [47] to understand the underlying issues. The insights provided by this model revealed that the models disproportionately focused on the images’ background rather than on the toy rockets themselves, as shown in Figure 3. This misdirection was the primary cause of the poor performance. By cropping the images to remove the background, we ensured that the models concentrated on the relevant features of the toy rockets. This adjustment improved the model performance, leading to more accurate anomaly detection. The final cropped images are of two sizes, determined by the cycle states: 200 pixels by 70 pixels for cycle state 4 and 400 pixels by 205 pixels for cycle state 9.
Figure 3: Score-CAM visualization highlighting the model’s focus on background elements emphasizing the need for cropping and object detection to improve accuracy by isolating the toy rocket.
@@ III-D Text Guided Zero-Shot Object Detection
We then incorporated a zero-shot object detection model to enhance the accuracy and robustness of our anomaly detection pipeline targeting the detection of toy rockets. Specifically, we employed OWL-ViT [48], a model that leverages multi-modal representations to perform open-vocabulary detection. This approach allows for the detection of objects based on free-text prompts, facilitating a flexible and powerful detection mechanism. OWL-ViT integrates CLIP (Contrastive Language-Image Pretraining) with lightweight object classification and localization heads. This integration enables the model to handle open-vocabulary detection by embedding free-text queries through CLIP’s text encoder, which are then utilized as inputs for the object classification and localization heads. The Vision Transformer (ViT) processes image patches as inputs, associating them with their corresponding textual descriptions.
III-E Training a Custom Object Detection Model with YOLOv9
To further enhance our anomaly detection pipeline for toy rockets, we trained a custom object detection model using YOLOv9 [19]. While zero-shot models like OWL-ViT offer significant flexibility and adaptability, a custom-trained model provides specific advantages such as:
  • By training YOLOv9 specifically on our dataset of toy rockets, we can fine-tune the model parameters to achieve higher accuracy and precision for this specific task, compared to a more general-purpose zero-shot model.
  • YOLOv9 is designed for high-speed inference, making it ideal for real-time applications in a manufacturing pipeline. In contrast, zero-shot models like OWL-ViT, while flexible, may have higher computational requirements and longer inference times.
@@ IV Problem Formulation
Consider a dataset $\mathcal{D}$ comprising images captured at various stages of the toy rocket manufacturing process. The manufacturing process is divided into cycles, each consisting of 21 distinct states. Let $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{n}\}$ denote the set of manufacturing cycles, where each cycle $C_{j}$ for $j=1,2,\ldots,n$ represents the complete assembly of one toy rocket.
Each cycle $C_{j}$ is divided into 21 states, represented by the set $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{21}\}$. For our anomaly detection task, we focus on images corresponding to cycle states $S_{4}$ and $S_{9}$, where rocket parts are most likely to be visible.
Let =$\mathcal{I}=\{I_{j,s}\mid j=1,2,\ldots,n,\,s\in\{4,9\}\}$ denote the set of images captured during these specific cycle states. Each image $I_{j,s}$ is associated with a label $y_{j,s}\in\mathcal{L}$, where =$\mathcal{L}=\{\text{normal},\text{anomaly}_{1},\text{anomaly}_{2},\ldots,\text{anomaly}_{k}\}$. Additionally, let $B_{j,s}=(x_{\min},y_{\min},x_{\max},y_{\max})$ denote the bounding box for image $I_{j,s}$, obtained via object detection techniques to focus on the relevant part of the image.
We define the dataset as:
$\mathcal{D}=\{(I_{j,s},y_{j,s},B_{j,s})\mid j=1,2,\ldots,n,\,s\in\{4,9\}\}$

### v2

@@ Abstract
Anomaly detection in manufacturing pipelines remains a critical challenge, intensified by the complexity and variability of industrial environments. This paper introduces AssemAI, an interpretable image-based anomaly detection system tailored for smart manufacturing pipelines. Utilizing a curated image dataset from an industry-focused rocket assembly pipeline, we address the challenge of imbalanced image data and demonstrate the importance of image-based methods in anomaly detection. Our primary contributions include deriving an image dataset, fine-tuning an object detection model YOLO-FF, and implementing a custom anomaly detection model for assembly pipelines. The proposed approach leverages domain knowledge in data preparation, model development and reasoning. We implement several anomaly detection models on the derived image dataset, including a Convolutional Neural Network, Vision Transformer (ViT), and pre-trained versions of these models. Additionally, we incorporate explainability techniques at both user and model levels, utilizing ontology for user-level explanations and SCORE-CAM for in-depth feature and model analysis. Finally, the best-performing anomaly detection model and YOLO-FF are deployed in a real-time setting. Our results include ablation studies on the baselines and a comprehensive evaluation of the proposed system. This work highlights the broader impact of advanced image-based anomaly detection in enhancing the reliability and efficiency of smart manufacturing processes. The image dataset, codes to reproduce the results and additional experiments are available at https://github.com/renjithk4/AssemAI.
@@
Smart Manufacturing, Object Detection, Image Processing, Anomaly Detection, Interpretability
@@ I Introduction
Anomaly detection is an essential methodology in manufacturing to identify deviations from normal production processes, which can indicate potential issues such as equipment failures, defects, or inefficiencies. Anomaly detection helps ensure product quality, reduce downtime, and optimize production processes [3] [4]. By monitoring various process parameters and product characteristics, manufacturers can detect anomalies early and prevent costly production disruptions [5].
@@
Image data are increasingly being leveraged in manufacturing systems due to advancements in computer vision and the availability of high-resolution cameras. In modern manufacturing facilities, image data are utilized for a wide range of applications, including quality inspection [6, 7], predictive maintenance [8, 9], process monitoring [10] and process optimization [11]. The integration of image data allows for real-time monitoring and anomaly detection, significantly improving the ability to identify defects and streamline production processes. Specifically, image-based anomaly detection has gained prominence, leveraging visual data to identify defects or irregularities in products. Automated visual inspection systems use image data to detect surface defects, assembly errors and deviations from design specifications[12], providing real-time feedback essential for maintaining high production standards and minimizing defects[4].
@@
Despite the advancements in image-based anomaly detection, several challenges persist as follows: (i) In visual data, the distinctions between normal and anomalous conditions are often subtle, making it difficult for the model to differentiate these anomalies from typical image variations (ii) The need for large and diverse datasets for training can be resource-intensive, both in terms of data collection and annotation [5, 13]. (iii) Generalization issues that arise when a model trained in one manufacturing environment fails to adapt to others due to differences in lighting, camera angles, or product types [14]. (iv) Difficulty in interpreting detection results, hindering the ability to provide actionable insights for process improvement [15]. To address these challenges, this work presents the following contributions:
  1. 1.
    A new image dataset for anomaly detection in manufacturing, derived using a zero-shot object detection model OWL-ViT.
  2. 2.
@@
    A novel anomaly detection model that leverages the pre-trained EfficientNet architecture [16], fine-tuned on the derived dataset for enhanced anomaly detection in assembly pipelines.
  3. 3.
@@
    Integration of user-level explainability using process ontology and model-level explainability using SCORE-CAM for salient feature and model analysis,
Figure 1: Overall Architecture of AssemAI. The figure illustrates the AssemAI pipeline, beginning with dataset preparation, which includes filtering and cropping images. Next, object detection is performed using zero-shot detection and a fine-tuned model (YOLO-FF). This is followed by anomaly detection using CNN, Custom-ViT, Pre-trained-ViT, and EfficientNet. The detection output is explained using SCORE-CAM for model-level explanations and process ontology for user-level explanations.
@@ II-A Object Detection Models
Object detection has advanced significantly, with various models enhancing both accuracy and efficiency. Girshick et al. [17] introduced R-CNN, which applies CNNs to region proposals for object classification and bounding box refinement. Faster R-CNN [18] built on R-CNN by integrating a Region Proposal Network (RPN) for faster detection. The YOLO models, from YOLOv1 to YOLOv9, transformed object detection by formulating it as a single regression problem to predict bounding boxes and class probabilities directly [19]. ShuffleNet [20] and MobileNet [21] are lightweight models designed for efficiency on mobile devices. SqueezeNet [22] aimed to reduce model size while maintaining accuracy with its Fire module. Lastly, the Swin Transformer [23] introduced hierarchical transformers with shifted windows, setting new benchmarks for object detection and vision tasks. Despite these advancements, limitations such as the need for extensive labeled data and computational resources persist. Our work addresses these limitations by focusing on lightweight and efficient models tailored for anomaly detection in manufacturing settings, improving both detection speed and accuracy while maintaining interpretability.
@@ II-B Zero-Shot Object Detection
Zero-shot object detection (ZSD) recognizes objects without labeled training data but faces challenges like semantic noise and class imbalance. Foundational methods by Bansal et al. [24] and Rahman et al. [25] struggled with noise. Gupta et al. [26] and Zheng et al. [27] improved ZSD with symmetric mapping and cascade stages, but issues persisted. Hayat et al. [28] and Li et al. [29] used ResNet, KNN, and contextual information but faced noise and ambiguity. Zhu et al. [30] and Hayat et al. [31] advanced ZSD with Don’t Even Look Once (DELO) model and Generative-ZSD. Liu et al. [32] and Li et al. [33] proposed contrastive learning and a semantics-aware framework. Our work focuses on specialized domains for precise anomaly detection.
@@ II-C Anomaly detection
Anomaly detection has seen significant advancements across diverse domains like networking [34], smart agriculture [35], healthcare [36], manufacturing [37, 38, 39, 40, 41, 42, 43]. Haselmann et al. [44] present an unsupervised one-class learning method using a deep CNN for surface inspection, outperforming other methods on decorated plastic parts. Xie et al. [37] propose a uniform benchmark for assessing image anomaly detection (IAD) algorithms in industrial settings. Maggipinto et al. [38] use convolutional autoencoders for monitoring semiconductor manufacturing, enhancing effectiveness and scalability. Jiang et al. [39] introduce YOLOv3 for balanced datasets and Fast-AnoGAN for unbalanced datasets in industrial production. Tan et al. [40] use an encoder-decoder for anomaly detection in sequential sensor data. Kim et al. [41] propose a self-supervised method using Gramian angular field and StyleGAN for time-series data. Bougaham et al. [42] demonstrate a three-step deep learning approach for Printed Circuit Board Assembly (PCBA) images, achieving high accuracy. Despite advancements, publicly available datasets for assembly processes are scarce, and existing methods lack interpretability and explainability for domain-specific insights.
@@ III Proposed Methodology
Figure 1 shows the overall architecture of the proposed method. The AssemAI pipeline begins with dataset preparation, where images are filtered and cropped to focus on relevant features. We then use a zero-shot object detection model to get the region of interest for the rocket in the image, deriving a dataset with masked regions of interest and bounding box details in a CSV file. A YOLO model is then fine-tuned on these images and bounding boxes. Subsequently, this derived dataset is used for anomaly detection experiments with several architectures, including CNN, Custom ViT, Pre-trained ViT, and EfficientNet.
@@
The detection output is then explained using SCORE-CAM for model-level explanations and integrated with process ontology for user-level explanations, providing technical and domain-specific insights.
III-A Future Factories Multimodal Dataset
We use the Future Factories (FF) dataset [45] curated and publicly released by the Future Factories lab at the McNair Aerospace Research Center, University of South Carolina. The dataset consists of measurements from a simulation of a rocket assembly pipeline, which adheres to industrial standards in deploying actuators, control mechanisms and transducers [^1] Refer to the images available in the GitHub repository for further details.. The dataset comprises two versions; analog and multimodal dataset, for which we use the images included in the multimodal dataset in our study. Table I shows the statistics of the original images in the FF multimodal dataset.
III-B Image filtering
The rocket assembly process at the FF lab is divided into 21 distinct cycle states. Information about these cell cycle states is not directly available in the multimodal dataset and had to be extracted from the analog dataset using a mapping function provided by domain experts. By calculating the Structural Similarity Index (SSIM) between the normal and anomalous images as shown in Figure 2, we observed that the images are structurally very similar across most cycle states. This indicates that the differences between the images are subtle and localized, which suggests that further analysis is needed to filter and figure out the region of interest. Among these cell cycle states, the rocket and its parts are visible only in two specific cycle states due to the spatial location of robots and other machinery and the location of the cameras and camera angles. We focus on filtering cycle four and a section of cycle nine based on domain knowledge and observational insights for enhanced image understanding. Table II and III show the filtered dataset and its anomaly statistics, respectively.
Figure 2: Structural Similarity between Normal and Anomalous Images
@@ III-C Image Cropping
Our dataset comprises images captured at various stages of the manufacturing cycle. In each cycle, the rocket’s position remains consistent within each respective state. Using this domain knowledge, we defined a cropping strategy by defining a bounding box for each cycle state to remove the background from all images. This consistent positioning of the rockets across cycles enables us to crop the images accurately, ensuring that only the relevant parts of the images are retained.
The suboptimal results obtained with uncropped images necessitated cropping of images. Specifically, our models performed poorly, with significant misclassification rates. We employed the explainability model SCORE-CAM [46] to understand the underlying issues. The insights provided by SCORE-CAM revealed that the models prioritized images’ backgrounds rather than the rockets themselves, as shown in Figure 3. Therefore, images were cropped to remove the backgroun. This image transformation improved the model performance, leading to better performance. As the rocket is the most important feature, the images are cropped accordingly, which resulted in a 200x70 pixel image for cycle state four and 400x205 pixels for cycle state nine.
Figure 3: Score-CAM visualization shows the model mistakenly focusing on background elements, highlighted in red as the most important regions, rather than the assembly pipeline marked in blue. This emphasizes the need for cropping and object detection to improve accuracy by isolating the relevant parts of the image.
@@ III-D Text Guided Zero-Shot Object Detection
We then incorporate a zero-shot object detection model to enhance the accuracy and robustness of our anomaly detection pipeline, which targets the detection of rockets and their parts. Specifically, we employ OWL-ViT [47], a model that leverages multimodal representations to perform open-vocabulary detection. This approach allows for the detection of objects based on free-text prompts, facilitating a flexible and powerful detection mechanism. OWL-ViT integrates CLIP (Contrastive Language-Image Pretraining) with lightweight object classification and localization heads. This integration enables the model to handle open-vocabulary detection by embedding free-text queries through CLIP’s text encoder. These are then utilized as inputs for the object classification and localization heads. The ViT processes image patches as inputs, associating them with their corresponding textual descriptions.
III-E Fine-tuning Object Detection Model with YOLOv9
To further enhance the object detection pipeline for rocket assembly, we fine-tune a YOLOv9 object detection model [19] using the dataset derived from OWL-ViT, which we refer to as YOLO-FF. While zero-shot models like OWL-ViT offer significant flexibility and adaptability, a fine-tuned model provides specific advantages as follows:
  • Fine-tuning YOLOv9 using the derived dataset ensures the model parameters are optimized for the specific task, facilitating better accuracy and precision.
  • YOLOv9 is optimized for high-speed inference, making it ideal for real-time applications, whereas models like OWL-ViT typically require more computation and have longer inference times.
@@ IV Problem Formulation
Consider a dataset $\mathcal{D}$ comprising images captured at various stages of the rocket assembly process. The manufacturing process is divided into cycles, each consisting of 21 distinct states. Let $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{n}\}$ denote the set of manufacturing cycles, where each cycle $C_{j}$ for $j=1,2,\ldots,n$ represents the complete assembly of one rocket. Each cycle $C_{j}$ is divided into 21 states, represented by the set $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{21}\}$. For our anomaly detection task, we focus on images corresponding to cycle states $S_{4}$ and $S_{9}$, where rocket parts are most likely to be visible. Let =$\mathcal{I}=\{I_{j,s}\mid j=1,2,\ldots,n,\,s\in\{4,9\}\}$ denote the set of images captured during these specific cycle states. Each image $I_{j,s}$ is associated with a label $y_{j,s}\in\mathcal{L}$, where =$\mathcal{L}=\{\text{normal},\text{anomaly}_{1},\text{anomaly}_{2},\ldots,\text{anomaly}_{k}\}$. Additionally, let $B_{j,s}=(x_{\min},y_{\min},x_{\max},y_{\max})$ denote the bounding box for image $I_{j,s}$, obtained via object detection techniques to focus on the relevant part of the image. We define the dataset as in equation 1.
$\mathcal{D}=\{(I_{j,s},y_{j,s},B_{j,s})\mid j=1,2,\ldots,n,\,s\in\{4,9\}\}$ (1)